# Computational Complexity of Physical Information Sufficiency

### v3

@@ 6 Tractable Special Cases: When You Can Solve It
We distinguish the encodings of Section 2.4. The tractability results below state the model assumption explicitly. Structural insight: hardness dissolves exactly when the full decision boundary $s\mapsto\text{Opt}(s)$ is recoverable in polynomial time from the input representation; the three cases below instantiate this single principle. Concretely, each tractable regime corresponds to a specific structural insight (explicit boundary exposure, additive separability, or tree factorization) that removes the hardness witness; this supports reading the general-case hardness as missing structural access in the current representation rather than as an intrinsic semantic barrier. In physical terms, these are the regimes where required structure is directly observable or factorized so that verification work scales with exposed local structure rather than hidden global combinations on encoded physical instances.[C:T6.1,2.37;P2.34;R:E,TA,AR] (L: CC69,CC70,CC71,CC72,CT6,CT4.)
@@
    Explicit-state encoding: The input contains the full utility table over $A\times S$. SUFFICIENCY-CHECK runs in $O(|S|^{2}|A|)$ time; if $|A|$ is constant, $O(|S|^{2})$.
@@
    Separable utility (any encoding): $U(a,s)=f(a)+g(s)$.
@@
    Tree-structured utility with explicit local factors (succinct structured encoding): There exists a rooted tree on coordinates and local functions $u_{i}$ such that
@@
    with the root term depending only on $(a,s_{\mathrm{root}})$ and all $u_{i}$ given explicitly as part of the input.
(L: CC69, CC70, CC72, CC71.)
6.1 Explicit-State Encoding
@@ 6.2 Separable Utility
If $U(a,s)=f(a)+g(s)$, then:
$\text{Opt}(s)=\arg\max_{a\in A}[f(a)+g(s)]=\arg\max_{a\in A}f(a)$
The optimal action is independent of the state. Thus $I=\emptyset$ is always sufficient.  $\blacksquare$
6.3 Tree-Structured Utility
@@ 6.4 Practical Implications
Explicit-state encoding Small or fully enumerated state spaces
@@
Tree-structured utility Hierarchies, causal trees
@@
[C:P6.2;R:H=tractable-structured] Let $\mathcal{C}$ be a structure class for which SUFFICIENCY-CHECK is polynomial (Theorem 6.1). If membership in $\mathcal{C}$ is decidable in polynomial time, then the combined detect-then-check procedure is a sound, polynomial-time heuristic applicable to all future instances in $\mathcal{C}$. For each subcase of Theorem 6.1, structure detection is polynomial (under the declared representation assumptions). (L: CC53, CC7, CC55, CC74.)
@@
Proposition 6.2 removes the complexity-theoretic obstruction to integrity-preserving action on detectable tractable instances: integrity no longer forces abstention. Whether the agent acts still requires competence (Definition 3.9): budget and coverage must also be satisfied. An agent that detects structure class $\mathcal{C}$, applies the corresponding polynomial checker, and abstains when detection fails maintains integrity; it never claims sufficiency without verification. Mistaking the heuristic for the general solution, claiming polynomial-time competence on a coNP-hard problem, violates integrity (Proposition 3.61).
@@ 7 Regime-Conditional Corollaries
This section derives regime-typed engineering corollaries from the core complexity theorems. Theorem 7.2 maps configuration simplification to SUFFICIENCY-CHECK; Theorems 4.6, 4.7, and 5.1 then yield exact minimization consequences under [S] and [S+ETH]. The same derivation applies to physical configuration/control systems once they are encoded into the C1–C4 decision interface: “configuration parameter” becomes any declared physical control or observation coordinate, and the same sufficiency predicate governs behavior preservation. [C:T7.2,2.37;P2.34;R:AR] (L: CR1,CT6,CT4.)
Regime tags used below follow Section 2.5: [S], [S+ETH], [E], [S_bool]. Any prescription that requires exact minimization is constrained by these theorem-level bounds.[C:T4.6,4.7,5.1;R:S,S+ETH] Theorem 7.4 implies that persistent failure to isolate a minimal sufficient set is a boundary-characterization signal in the current model, not a universal irreducibility claim.[C:T7.4;R:S]
@@
For the objective “minimize verified total cost while preserving integrity,” persistent over-specification is admissible only in the attempted competence failure cell (Definition 3.59): after an attempted exact procedure, if exact irrelevance cannot be certified efficiently, integrity retains uncertified coordinates rather than excluding them.[C:P3.60;R:AR] When exact competence is available in the active regime (e.g., Theorem 6.1 and the exact-identifiability criterion), proven-irrelevant coordinates are removable; persistent over-specification is irrational for that same objective.[C:T6.1;P3.60;R:E,STR] Proposition 3.60 makes this explicit: in the integrity-preserving matrix, one cell is rational and three are irrational, so irrationality is the default verdict.[C:P3.60;R:AR] (L: IC18, IC25, IC7, CC71, S2P1.)
@@
All claims in this section are formal corollaries under the declared model assumptions.
@@
  • Therefore, hardness results imply a regime-conditional trilemma: abstain, weaken guarantees (heuristics/approximation), or change encoding/structural assumptions to recover competence.[C:T4.6,5.1;P3.60;R:S,S+ETH]
@@ 7.2 Cost Asymmetry Under ETH
Within [S+ETH], persistent over-specification is unresolved boundary characterization, not proof that all included parameters are intrinsically necessary. Conversely, when exact competence is available in the active regime, certifiably irrelevant coordinates are removable; persistence is irrational for the same cost-minimization objective.[C:T7.5;P3.60;R:S+ETH,TA] (L: CC12, HD13.)
@@
The eventual dominance step is mechanized in HD8: for fixed linear-overhead parameter $k$ and additive constant $C_{\text{under}}$ there is $n_{0}$ such that $k<2^{n}+C_{\text{under}}$ for all $n\geq n_{0}$. Therefore:
@@
[C:C7.6;R:H=succinct-hard] Assuming $\textsf{P}\neq\textsf{coNP}$, there exists no polynomial-time algorithm that:
@@ 7.3 Regime-Conditional Operational Corollaries
Theorems 7.4 and 7.5 yield the following conditional operational consequences:
1. Conservative retention under unresolved relevance. If irrelevance cannot be certified efficiently under the active regime, the admissible policy is to retain a conservative superset of parameters.[C:T7.4;R:S]
2. Heuristic selection as weakened-guarantee mode. Under [S+ETH], exact global minimization can be exponentially costly in the worst case (Theorem 7.5); methods such as AIC/BIC/cross-validation instantiate the “weaken guarantees” branch of Definition 3.9 [1, 56, 66, 25, 50].[C:T7.5;R:S+ETH]
3. Full-parameter inclusion as an $O(n)$ upper-bound strategy. Under [S+ETH], if exact minimization is unresolved, including all $n$ parameters is an $O(n)$ maintenance upper-bound policy that avoids false irrelevance claims.[C:T7.5;R:S+ETH]
4. Irrationality outside attempted-competence-failure conditions. If the active regime admits exact competence (tractable structural-access conditions or exact relevance identifiability), or if exact competence was never actually attempted, continued over-specification is irrational; hardness no longer justifies it for the stated objective.[C:T6.1;P3.60;R:TA] (L: CC71, IC6, S2P1.)
These corollaries are direct consequences of the hardness/tractability landscape: over-specification is an attempted-competence-failure policy, not a default optimum. Outside that cell, the admissible exits are to shift to tractable regimes from Theorem 6.1 or adopt explicit approximation commitments.[C:T6.1;P3.60;R:RG]
@@ 8 Dominance Theorems for Hardness Placement
Regime for this section: the mechanized Boolean-coordinate model [S_bool] plus the architecture cost model defined below. The centralization-vs-distribution framing follows established software-architecture criteria about modular decomposition and long-run maintenance tradeoffs [49, 10, 6, 60]. Although phrased in software terms, the formal objects are substrate-agnostic: the same hardness-placement and amortization equations apply to physical architectures (for example, centralized vs distributed sensing/actuation computation) once mapped into the same coordinate/cost model. [C:P2.34;T2.37,8.16,9.3,9.6;R:AR,LC] (L: CT6,CT4,HD1,HD4,HD3.)
@@ 8.1 Over-Specification as Diagnostic Signal
[C:C8.1;R:H=succinct-hard] In the mechanized Boolean-coordinate model, if a coordinate is relevant and omitted from a candidate set $I$, then $I$ is not sufficient. (L: S2P6)
@@
[C:C8.2;R:H=succinct-hard] In the mechanized Boolean-coordinate model, for any candidate set $I$:
@@
In this paper’s formal typing, persistent over-specification is admissible only under attempted competence failure: after an attempted exact procedure, exact relevance competence remains unavailable in the active regime, so integrity retains uncertified coordinates (Section 3; Section 7).[C:P3.60;R:AR] Once exact competence is available in the active regime (Corollaries 8.78.9 together with Corollary 8.2), certifiably irrelevant coordinates are removable; persistent over-specification is irrational for the same objective (verified total-cost minimization).[C:T6.1;P3.60;R:TA] (L: IC18, IC25, IC7, L15, L16, L17, S2P1.)

### v4

@@ 10 Tractable Special Cases: When You Can Solve It
Formally: this section gives sufficient structure/representation conditions for polynomial-time SUFFICIENCY-CHECK.
We distinguish the encodings of Section 2.4. The tractability results below state the model assumption explicitly. Structural insight: hardness dissolves exactly when the full decision boundary $s\mapsto\text{Opt}(s)$ is recoverable in polynomial time from the input representation; the six cases below instantiate this single principle. Concretely, each tractable regime corresponds to a specific structural insight (explicit boundary exposure, additive separability, tensor factorization, tree structure, bounded treewidth, or coordinate symmetry) that removes the hardness witness; this supports reading the general-case hardness as missing structural access in the current representation rather than as an intrinsic semantic barrier. In physical terms, these are the regimes where required structure is directly observable or factorized so that verification work scales with exposed local structure rather than hidden global combinations on encoded physical instances.[C:T10.1, 2.45;P2.42;R:E, TA, AR] (L: CC69, CC70, CC71, CC72, CC73, CT6, CT4.)
Informally: these cases are tractable because useful structure is visible.
@@
    Bounded actions (explicit-state encoding): The input contains the full utility table over $A\times S$. SUFFICIENCY-CHECK runs in $O(|S|^{2}|A|)$ time; if $|A|$ is constant, $O(|S|^{2})$.
@@
    Separable utility (rank 1): $U(a,s)=f(a)+g(s)$. The optimal action is state-independent, so $I=\emptyset$ is always sufficient. Complexity: $O(1)$ for sufficiency verification.
@@
    Low tensor rank: $U(a,s)=\sum_{r=1}^{R}w_{r}\cdot f_{r}(a)\cdot\prod_{i=1}^{n}g_{ri}(s_{i})$ for bounded $R$. Complexity: $O(|A|\cdot R\cdot n)$.
  4. 4.
    Tree-structured dependencies: Dependencies form a rooted tree:
@@
    with the root term depending only on $(a,s_{\mathrm{root}})$ and all $u_{i}$ given explicitly. Complexity: $O(n\cdot|A|\cdot k_{\max})$.
  5. 5.
    Bounded treewidth: Utility decomposes as pairwise interactions on an interaction graph of treewidth $w$. Complexity: $O(n\cdot k^{w+1})$ via CSP algorithms [12, 29].
  6. 6.
    Coordinate symmetry: State space is $S=[k]^{d}$ with utility invariant under coordinate permutations. The effective state space reduces from $k^{d}$ to $\binom{d+k-1}{k-1}$ orbit types. Complexity: $O\bigl(\binom{d+k-1}{k-1}^{2}\bigr)$, polynomial in $d$ for fixed $k$.
(L: CC70, CC71, CC72, CC73.)
10.1 Bounded Actions (Explicit-State)
(L: CC70.)
@@ 10.2 Separable Utility (Rank 1)
If $U(a,s)=f(a)+g(s)$, then $\text{Opt}(s)=\arg\max_{a\in A}f(a)$. The optimal action is independent of the state. Thus $I=\emptyset$ is always sufficient, and sufficiency verification is $O(1)$.  $\blacksquare$
10.3 Low Tensor Rank
(L: CC71.)
@@ 10.3 Low Tensor Rank
When utility has tensor rank $R$: $U(a,s)=\sum_{r=1}^{R}w_{r}f_{r}(a)\prod_{i=1}^{n}g_{ri}(s_{i})$, the problem reduces to factored tensor computation. For each action, the optimal value computation requires $O(R\cdot n)$ operations. Total complexity: $O(|A|\cdot R\cdot n)$, polynomial for bounded $R$. This reduction cites standard tensor contraction algorithms [40, 18].  $\blacksquare$
10.4 Tree-Structured Dependencies
(L: CC73.)
Proof of Part 4.
@@ 10.4 Tree-Structured Dependencies
10.5 Bounded Treewidth
(L: CC73, CC75.)
Proof of Part 5.
When utility decomposes as pairwise interactions $U(a,s)=\sum_{(i,j)}u_{ij}(a,s_{i},s_{j})$, the interaction graph has an edge between coordinates $i,j$ iff $u_{ij}$ is nontrivial. If this graph has treewidth $w$, SUFFICIENCY-CHECK reduces to a constraint satisfaction problem on a bounded-treewidth graph. By standard CSP algorithms [12, 29], this is solvable in $O(n\cdot k^{w+1})$ time.  $\blacksquare$
10.6 Coordinate Symmetry
(L: CC72.)
Proof of Part 6.
For state space $S=[k]^{d}$ with coordinate-permutation-invariant utility, two states have the same optimal actions iff they have the same orbit type (value histogram). The number of orbit types is $\binom{d+k-1}{k-1}$ by stars-and-bars. Sufficiency checking reduces to comparing optimal actions across orbit type pairs. Total complexity: $O\bigl(\binom{d+k-1}{k-1}^{2}\cdot|A|\bigr)$, polynomial in $d$ for fixed $k$.  $\blacksquare$
@@ 10.7 Practical Implications
Formally: this subsection states detect-then-check closure for the declared tractable classes.
@@
Bounded actions Small or fully enumerated state spaces
@@
Low tensor rank Factored decision trees, sparse interactions
Tree-structured Hierarchies, causal trees, DAGs
Bounded treewidth Local dependency graphs, grid-like structures
Coordinate symmetry Exchangeable particles, counting problems
@@
[C:P10.2;R:H=tractable-structured] Let $\mathcal{C}$ be a structure class for which SUFFICIENCY-CHECK is polynomial (Theorem 10.1). If membership in $\mathcal{C}$ is decidable in polynomial time, then the combined detect-then-check procedure is a sound, polynomial-time heuristic applicable to all future instances in $\mathcal{C}$. For each subcase of Theorem 10.1, structure detection is polynomial (under the declared representation assumptions). (L: CC53, CC7, CC55, CC74.)
@@
Proposition 10.2 removes the complexity-theoretic obstruction to integrity-preserving action on detectable tractable instances: integrity no longer forces abstention. Whether a declared controller executes the action still requires competence (Definition 3.9): budget and coverage must also be satisfied. A controller that detects structure class $\mathcal{C}$, applies the corresponding polynomial checker, and abstains when detection fails maintains integrity; it never claims sufficiency without verification. Mistaking the heuristic for the general solution, claiming polynomial-time competence on a coNP-hard problem, violates integrity (Proposition 3.61).
Informally: reusable heuristics are reliable only after the stated structure test; otherwise hard-case limits still apply.
@@ 11 Regime-Conditional Corollaries
Formally: this section derives operational corollaries by composing theorem-level hardness, tractability, and typing results under declared regimes.
This section derives regime-typed engineering corollaries from the full map: static/stochastic/sequential placements (Theorems 4.6, 6.7, 7.7) with strict hierarchy (Propositions 8.1, 8.2), and the static decision/access decomposition used operationally here. Theorem 11.2 maps configuration simplification to SUFFICIENCY-CHECK; Theorems 4.6, 4.7, and 9.1 then yield exact minimization consequences under [S] and [S+ETH]. The same derivation applies to physical configuration/control systems once they are encoded into the C1–C4 decision interface: “configuration parameter” becomes any declared physical control or observation coordinate, and the same sufficiency predicate governs behavior preservation. [C:T11.2, 4.6, 6.7, 7.7, 2.45;P2.42, 8.1, 8.2;R:AR, RG] (L: CR1, CT6, CT4.)
For deployed software, this remains physical at execution time: each configuration read/write and verification step is realized by electrical state transitions in hardware, so coordinate minimization and over-specification map directly to measured switching-energy load under the declared bit-to-energy lift. [C:P9.28, 2.36;R:AR, H=cost-growth] (L: CC76, DT16.)
Under the same discrete-time event law, more verification cycles mean more decision-event ticks and more bit operations; with positive per-bit conversion, this induces additional thermodynamically necessary heat-transfer load. [C:P2.35, 2.36, 9.28;C9.32;R:AR, H=cost-growth] (L: DT13, DT16, CV17, CV18, CC76, CC77.)
Regime tags used below follow Section 2.5: [S], [S+ETH], [E], [S_bool]. Any prescription that requires exact minimization is constrained by these theorem-level bounds.[C:T4.6, 4.7, 9.1;R:S, S+ETH] Theorem 11.4 implies that persistent failure to isolate a minimal sufficient set is a boundary-characterization signal in the current model, not a universal irreducibility claim.[C:T11.4;R:S]
@@
For the objective “minimize verified total cost while preserving integrity,” persistent over-specification is admissible only in the attempted competence failure cell (Definition 3.59): after an attempted exact procedure, if exact irrelevance cannot be certified efficiently, integrity retains uncertified coordinates rather than excluding them.[C:P3.60;R:AR] When exact competence is available in the active regime (e.g., Theorem 10.1 and the exact-identifiability criterion), proven-irrelevant coordinates are removable; persistent over-specification is irrational for that same objective.[C:T10.1;P3.60;R:E, STR] Proposition 3.60 makes this explicit: in the integrity-preserving matrix, one cell is rational and three are irrational, so irrationality is the default verdict.[C:P3.60;R:AR] (L: IC18, IC25, IC7, CC71, S2P1.)
@@
All claims in this section are formal corollaries under the stated model contract.
@@
  • Therefore, hardness results imply a regime-conditional trilemma: abstain, weaken guarantees (heuristics/approximation), or change encoding/structural assumptions to recover competence.[C:T4.6, 9.1;P3.60;R:S, S+ETH]
Informally: these recommendations come directly from proved results under stated regimes.
@@ 11.1 Configuration Simplification is SUFFICIENCY-CHECK
Formally: this subsection proves configuration-preservation queries instantiate SUFFICIENCY-CHECK under C1–C4.
@@
Informally: if a minimal relevant set cannot be certified, that indicates an unresolved model boundary, not proof that every retained parameter is necessary.
@@ 11.2 Cost Asymmetry Under ETH
Formally: this subsection proves asymptotic finding-vs-maintenance asymmetry under succinct ETH and the declared cost model.
@@
Within [S+ETH], persistent over-specification is unresolved boundary characterization, not proof that all included parameters are intrinsically necessary. Conversely, when exact competence is available in the active regime, certifiably irrelevant coordinates are removable; persistence is irrational for the same cost-minimization objective.[C:T11.5;P3.60;R:S+ETH, TA] (L: CC12, HD14.)
@@
The eventual dominance step is mechanized in HD14: for fixed linear-overhead parameter $k$ and additive constant $C_{\text{under}}$ there is $n_{0}$ such that $k<2^{n}+C_{\text{under}}$ for all $n\geq n_{0}$. Therefore:
@@
[C:C11.6;R:H=succinct-hard] Assuming $\textsf{P}\neq\textsf{coNP}$, there exists no polynomial-time algorithm that:
@@
Informally: at scale, exact minimization can cost more than linear maintenance in worst cases, while structured small-parameter families may still be tractable.
@@ 11.3 Regime-Conditional Operational Corollaries
Formally: this subsection instantiates admissible policy outcomes from prior theorems and integrity-competence typing.
Theorems 11.4 and 11.5 yield the following conditional operational consequences:
1. Conservative retention under unresolved relevance. If irrelevance cannot be certified efficiently under the active regime, the admissible policy is to retain a conservative superset of parameters.[C:T11.4;R:S]
2. Heuristic selection as weakened-guarantee mode. Under [S+ETH], exact global minimization can be exponentially costly in the worst case (Theorem 11.5); methods such as AIC/BIC/cross-validation instantiate the “weaken guarantees” branch of Definition 3.9 [1, 75, 86, 33, 65].[C:T11.5;R:S+ETH]
3. Full-parameter inclusion as an $O(n)$ upper-bound strategy. Under [S+ETH], if exact minimization is unresolved, including all $n$ parameters is an $O(n)$ maintenance upper-bound policy that avoids false irrelevance claims.[C:T11.5;R:S+ETH]
4. Irrationality outside attempted-competence-failure conditions. If the active regime admits exact competence (tractable structural-access conditions or exact relevance identifiability), or if exact competence was never actually attempted, continued over-specification is irrational; hardness no longer justifies it for the stated objective.[C:T10.1;P3.60;R:TA] (L: CC71, IC6, S2P1.)
These corollaries are direct consequences of the hardness/tractability landscape: over-specification is an attempted-competence-failure policy, not a default optimum. Outside that cell, the admissible exits are to shift to tractable regimes from Theorem 10.1 or adopt explicit approximation commitments.[C:T10.1;P3.60;R:RG]
Informally: policy choices are threefold: abstain, weaken guarantees, or change assumptions to recover competence.
@@ 12 Dominance Theorems for Hardness Placement
Regime for this section: the mechanized Boolean-coordinate model [S_bool] plus the architecture cost model defined below. The centralization-vs-distribution framing follows established software-architecture criteria about modular decomposition and long-run maintenance tradeoffs [63, 14, 6, 79]. Although phrased in software terms, the formal objects are substrate-agnostic: the same hardness-placement and amortization equations apply to physical architectures (for example, centralized vs distributed sensing/actuation computation) once mapped into the same coordinate/cost model. [C:P2.42;T2.45, 12.19, 13.3, 13.6;R:AR, LC] (L: CT6, CT4, HD1, HD5, HD4.)
In electrical implementations, the same architecture equations are energy equations: each distributed update path induces switching events and Joule expenditure, while centralized one-time structure pays setup cost and reduces repeated per-site electrical work. Under the declared thermodynamic lift, this is tracked as a typed transfer from bit-operation load to physical energy load. [C:P9.28;C9.32;R:AR, H=cost-growth] (L: CC76, CC77.)
Equivalently, repeated distributed cycles create a thermal duty: more update cycles imply more bit operations, which imply more unavoidable dissipated heat and therefore larger heat-transfer/removal requirements in the same substrate class. [C:P2.35, 2.36, 9.28;R:AR, H=cost-growth] (L: DT13, DT16, CV17, CV18, CC76.)
@@ 12.1 Over-Specification as Diagnostic Signal
[C:C12.1;R:H=succinct-hard] In the mechanized Boolean-coordinate model, if a coordinate is relevant and omitted from a candidate set $I$, then $I$ is not sufficient. (L: S2P6)
@@
[C:C12.2;R:H=succinct-hard] In the mechanized Boolean-coordinate model, for any candidate set $I$:
@@
In this paper’s formal typing, persistent over-specification is admissible only under attempted competence failure: after an attempted exact procedure, exact relevance competence remains unavailable in the active regime, so integrity retains uncertified coordinates (Section 3; Section 11).[C:P3.60;R:AR] Once exact competence is available in the active regime (Corollaries 12.712.12 together with Corollary 12.2), certifiably irrelevant coordinates are removable; persistent over-specification is irrational for the same objective (verified total-cost minimization).[C:T10.1;P3.60;R:TA] (L: IC18, IC25, IC7, CC81, CC82, CC84, CC71, CC73, CC72, S2P1.)